the resistance(R) of a wire is directly proportionnal to the lenght(L) and inversely proportional to the diameter(D) since a wire to 3 metres long with a diameter of o,25 centimetre has a resistance of 12 OHMS, calulate the length of a wire diameter of 0,3 centimetres and a resistance of 20 OHMS
Answer by stanbon(75887)

the resistance(R) of a wire is directly proportionnal to the lenght(L) and inversely proportional to the diameter(D) since a wire to 3 metres long with a diameter of o,25 centimetre has a resistance of 12 OHMS, calulate the length of a wire diameter of 0,3 centimetres and a resistance of 20 OHMS
--------------
R = k*L/D
---------
Find k:
12 = k*3/0.25
3 = 3k
k = 1
----------
EQUATION:
R = L/D
20 = L/0.3
L = 6 meters
==================
Cheers,
Stan H.